40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 26 17:50:22 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
    Reply  Wed Jan 27 18:19:45 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement 
       Reply  Thu Jan 28 00:29:30 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement PLL_OLG.pdfFreq_noise_comparison.pdf
          Reply  Mon Feb 1 18:11:25 2016, gautam, Update, Green Locking, Lightwave frequency noise measurement Free_running_frequency_noise_comparison.pdf
       Reply  Fri Jan 29 14:43:47 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.pdf
          Reply  Fri Jan 29 16:55:27 2016, rana, Update, Green Locking, Innolight laser is 10 years old 
             Reply  Sat Jan 30 00:12:22 2016, gautam, Update, Green Locking, Innolight laser is 10 years old RIN_comparison.pdf
                Reply  Sat Jan 30 09:56:24 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                   Reply  Mon Feb 1 15:16:28 2016, gautam, Update, Green Locking, Innolight laser is 10 years old 
                      Reply  Mon Feb 1 15:43:18 2016, Koji, Update, Green Locking, Innolight laser is 10 years old 
                         Reply  Fri Feb 5 15:02:13 2016, gautam, Update, Green Locking, X-end NE cable 
          Reply  Fri Feb 12 11:10:49 2016, Steve, Update, Green Locking, Innolight laser is 10 years old inno1W.jpg
          Reply  Mon Mar 21 14:29:32 2016, Steve, Update, Calibration-Repair, 1W Innolight laser repair diagnoses 
             Reply  Thu Mar 24 07:56:09 2016, Steve, Update, Calibration-Repair, NO Noise Eater for 1W Innolight 
                Reply  Mon Apr 11 17:03:41 2016, Steve, Update, Calibration-Repair, 1W Innolight repair completed 
Message ID: 11951     Entry time: Tue Jan 26 17:50:22 2016     Reply to this: 11953
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Lightwave frequency noise measurement 

I attempted to measure the frequency noise of the extra Lightwave NPRO we have that is currently sitting on the PSL table. I did the following:

  1. Turn the Lightwave NPRO back on.
  2. Disable MC autolocker and close the PSL shutter.
  3. Checked the alignment of the pick off from the PSL beam and the beam from the Lightwave NPRO onto the PDA10CF. These seemed okay, and I didn't really have to tweak any of the steering optics. I was getting a DC signal level of ~7V (the PD should drive a 1Mohm load up to 10V so it wasn't saturated).
  4. Swept the crystal temperature on the Lightwave using the dial on the front panel of the controller. I found beatnotes at 48.1831 degrees and 45.3002 degrees. However, the amplitude of the beatnote was pretty small (approx. -40dBm on the Agilent NA). I tried playing around with the beam alignment and laser power on the Lightwave NPRO to see if I could increase the beatnote amplitude, but was unsuccessful - turning up the laser power (from the nominal level of 55mW as per the front panel display) caused the PD to saturate at 10V, while as far as I could tell, the alignment of the two beams onto the PD is reasonably good. This seems inconsistent with the numbers Koji has reported in this elog, where he was able to get a beatnote of ~1Vpp for a DC of 2.5 V. 
  5. I tried locking the PLL (in roughly the same configuration as reported here) with this small amplitude beatnote but was unsuccessful. 

I've turned the Lightwave NPRO back to standby for now, in anticipation of further trials later today. I've also restored the IMC. 

ELOG V3.1.3-